### v2

# The affine closure of cotangent bundles of horospherical spaces

Abstract.
For a smooth quasi-affine variety $X$, the affine closure $\overline{T^{*}X}\coloneqq\operatorname{Spec}(k[T^{*}X])$ contains $T^{*}X$ as an open subset, and its smooth locus carries a symplectic structure. A natural question is whether $\overline{T^{*}X}$ itself is a symplectic variety. A notable example is the conjecture of Ginzburg and Kazhdan, which predicts that $\overline{T^{*}(G/U)}$ is symplectic for a maximal unipotent subgroup $U$ in a reductive linear algebraic group $G$. This conjecture was recently proved by Gannon using representation-theoretic methods.
In this paper, we provide a new geometric approach to this conjecture. Our method allows us to prove a more general result: $\overline{T^{*}(G/H)}$ is symplectic for any horospherical subgroup $H$ in $G$ such that $G/H$ is quasi-affine. In particular, this implies that the affine closure $\overline{T^{*}(G/[P,P])}$ is a symplectic variety for any parabolic subgroup $P$ in $G$.
Key words and phrases:
symplectic singularities, nilpotent orbits, basic affine spaces, horospherical varieties, cotangent bundles
2020 Mathematics Subject Classification:
14J42,22E10,14B05,14M17,14E30
1. Introduction
Throughout this paper, we work over an algebraically closed field $k$ of characteristic zero.
1.1. Affine closures of cotangent bundles
Let $X$ be an irreducible normal variety, and let $k[X]$ denote the $k$-algebra of regular functions on $X$. Since $k[X]$ is an integrally closed domain, we can define the affinization (or affine closure) of $X$ as the affine normal scheme $\overline{X}\coloneqq\operatorname{Spec}k[X]$. In this paper, we primarily investigate the affine closure of the cotangent bundle $T^{*}X$ of a smooth variety $X$.
The algebra $k[T^{*}X]$ is isomorphic to the algebra of symmetric tensors
$S(X)\coloneqq\bigoplus_{p\geq 0}H^{0}(X,{\sf S}^{p}T_{X})=\bigoplus_{p\geq 0}H^{0}(\mathbb{P}T_{X},\mathscr{O}_{\mathbb{P}T_{X}}(p)),$
where $\mathbb{P}E\coloneqq\operatorname{Proj}_{X}{\sf S}^{\bullet}E$ denotes the Grothendieck projectivization of a vector bundle $E$ over $X$. The symplectic structure on $T^{*}X$ induces a Poisson structure on $S(X)$ known as the Schouten–Nijenhuis bracket. Consequently, the affine closure $\overline{T^{*}X}$ (denoted by $\mathcal{Z}_{X}$ in [FL24]) forms a normal integral affine Poisson scheme. This construction yields a natural dominant affinization morphism
$\varphi_{X}:T^{*}X\to\overline{T^{*}X},$
which is generally non-surjective. Note that in general, $S(X)$ is not finitely generated, even for projective manifolds (see Section 3.2).
Despite its straightforward definition, the affine scheme $\overline{T^{*}X}$ and the morphism $\varphi_{X}$ have received limited attention in the literature (cf. [BL24, FL24]). In this context, the following fundamental questions naturally arise:
Question 1.1.
  1. (1.1.1)
    When is $S(X)$ finitely generated (over $k$)?
  2. (1.1.2)
    What kind of singularities can $\overline{T^{*}X}$ have?
Since $T^{*}X$ has a natural symplectic structure, it is natural to consider symplectic singularities as introduced by Beauville [Bea00]. Recall that a normal variety $W$ is symplectic if there exists a symplectic form $\omega$ on the smooth locus $W_{\text{reg}}$ of $W$ such that for any resolution $Z\to W$, the $2$-form $\omega$ extends to a regular $2$-form on $Z$. Moreover, a resolution $Z\to W$ is called a symplectic resolution if $\omega$ extends to a symplectic form on $Z$.
If $X$ is affine, then $T^{*}X$ is also affine, and thus $\varphi_{X}$ becomes an identity, implying that $\overline{T^{*}X}=T^{*}X$ is again a smooth symplectic variety. When $X$ is projective, it has been proven in [FL24] that $\varphi_{X}$ is birational if and only if $T_{X}$ is big. In this case, the affine closure $\overline{T^{*}X}$ is a symplectic variety if and only if $\mathbb{P}T_{X}$ is of Fano type. A typical example is when $X$ is a projective rational homogeneous space; in this case, $\varphi_{X}$ is the Stein factorization of the moment map, and $\overline{T^{*}X}$ is a finite cover of a nilpotent orbit closure in $\mathfrak{g}$.
1.2. Braverman–Kazhdan and horospherical spaces
In this paper, we investigate Question 1.1 for certain quasi-affine homogeneous spaces, motivated by the study of the basic affine space $G/U$, where $G$ is a connected reductive group and $U$ is the unipotent radical of a Borel subgroup $B$. The significance of $G/U$ and its cotangent bundle is highlighted in the following passage from a paper by J. Bernstein, I. M. Gelfand, and S. Gelfand [BGG75, p. 21]:
Experience in representation theory suggests that, for many problems in representation theory, the solution results from a careful study of the base affine space. In particular, the structure of the ring of differential operators on $G/U$ seems to be closely connected with the representations of the real forms of $G$.
In this direction, the finite generation of $S(G/U)$ was proved for $G=\operatorname{SL}_{n}$ in [DKS13] and was proved for general reductive groups in [GR15]. Moreover, V. Ginzburg and D. Kazhdan conjectured in [GK22, Conjecture 1.3.6] that the affine closure $\overline{T^{*}(G/U)}$ is a symplectic variety. This conjecture was verified by B. Jia in [Jia21] for $G=\mathrm{SL}_{n}$ and by T. Gannon in [Gan24] for general $G$. The key idea of their proofs is to use representation theory to show that the codimension of the singular locus is at least four, which in fact establishes the stronger result that $\overline{T^{*}(G/U)}$ has terminal symplectic singularities.
Since $U=[B,B]$, it is natural to ask whether the same holds for the Braverman–Kazhdan space $G/[P,P]$, where $P$ is a parabolic subgroup of $G$. Recall that a closed subgroup $H\subset G$ is called horospherical if its normalizer $N_{G}(H)$ is parabolic and the quotient $N_{G}(H)/H$ is a torus. It is known that $[P,P]$ is horospherical. Our first result confirms a generalization of the Ginzburg–Kazhdan conjecture for quasi-affine horospherical homogeneous spaces (see Theorem 4.5).
Theorem 1.2.
Let $G$ be a connected semi-simple group, and let $H\subset G$ be a horospherical subgroup such that $G/H$ is quasi-affine. Then $S(G/H)$ is finitely generated, and $\overline{T^{*}(G/H)}$ is symplectic.
As a corollary, we obtain the following result:
Corollary 1.3.
Let $G$ be a connected reductive group, and let $P$ be a parabolic subgroup of $G$.
  1. (1)
    Then $S(G/[P,P])$ is finitely generated, and $\overline{T^{*}(G/[P,P])}$ is a $\mathbb{Q}$-factorial symplectic variety.
  2. (2)
    If $[G,G]$ is simply connected, then $\overline{T^{*}(G/[P,P])}$ is factorial.
As a refinement of Question 1.1, we may consider the case where $X$ is a smooth quasi-affine variety so that $T^{*}X$ is also quasi-affine. It then follows from Proposition 3.2 that $\overline{T^{*}X}$ is symplectic if and only if $S(X)$ is finitely generated and $\overline{T^{*}X}$ has rational singularities. This leads to the following natural question:
Question 1.4.
Let $Z\subset\mathbb{P}^{n}$ be a smooth projective variety with $\widehat{Z}\subset\mathbb{A}^{n+1}$ its affine cone. Let $X=\widehat{Z}\setminus\{0\}$. When is $\overline{T^{*}X}$ a symplectic variety?
We will show in Proposition 4.7 that if $Z$ is a rational homogeneous space, then $\overline{T^{*}X}$ is a symplectic variety. In particular, this implies that $\overline{T^{*}\mathcal{O}_{\min}}$ is a symplectic variety for the minimal nilpotent orbit $\mathcal{O}_{\min}$ in a simple Lie algebra. It seems plausible that this property holds for all nilpotent orbits in a semi-simple Lie algebra.
In general, it is difficult to describe the affine closure $\overline{T^{*}X}$, even when it is symplectic. Surprisingly, it turns out that if $Z\subset\mathbb{P}^{n}$ is the variety of lines through a fixed point of an irreducible Hermitian symmetric space of rank at least two, then $\overline{T^{*}X}$ is the minimal nilpotent orbit closure in $\mathfrak{g}$, as demonstrated below.
Theorem 1.5.
Let $G$ be a connected simple group, and let $G/P$ be an irreducible Hermitian symmetric space of rank at least two. The nilradical $\mathfrak{u}_{P}\subset\mathfrak{p}$ is an irreducible $P$-module. Let $X\subset\mathfrak{u}_{P}$ be the orbit of a highest weight vector. Then, $\overline{T^{*}X}$ is isomorphic to the minimal nilpotent orbit closure $\overline{\mathcal{O}_{\text{min}}}\subset\mathfrak{g}$.
Our statement is slightly more general and covers certain associated varieties of unitary highest weight modules. For further details, see Theorem 4.9 and [LSS88, LS89]. It follows from Theorem 1.5 that if $\mathfrak{g}$ is not of type $G_{2}$, $F_{4}$, or $E_{8}$, then $\mathcal{O}_{\text{min}}$ contains a large open subset isomorphic to $T^{*}X$, which is somewhat surprising. Additionally, employing different methods, the closures of minimal nilpotent orbits in $\mathfrak{so}_{2n}$ ([Jia25]) and $\mathfrak{e}_{6}$ ([Jia25, GW25]) have also been realized as the affine closures of cotangent bundles (distinct from ours).
1.3. Parabolic basic affine spaces
Another natural generalization of the basic affine spaces is the parabolic basic affine spaces $X_{P}\coloneqq G/U_{P}$, where $U_{P}$ is the unipotent radical of a parabolic subgroup $P$ of a connected reductive group $G$. The affine closure $\overline{T^{*}(G/U)}$, called the universal implosion, was studied in the Dancer–Kirwan–Swann theory of hyperkähler implosions [DKS13] as well as in the context of $43d\ \mathcal{N}=$ 4 quiver gauge theory [BDG+22]. It is expected (cf. [DHK21, Section 8] and [BDG+22]) that $S(G/U_{P})$ is finitely generated and that $\overline{T^{*}(G/U_{P})}$, referred to as the partial implosion, is symplectic. This has been confirmed for $G=\operatorname{SL}_{n}$ or $\operatorname{GL}_{n}$ very recently by T. Gannon and B. Webster in [GW25].
After passing to an isogeny of simply connected type, the problem above can be reduced to the case where $G$ is (simply connected) semi-simple (see Lemma 5.1). Our final result reformulates the above expectation from the perspective of birational geometry. More precisely, let $\mathfrak{u}_{P}^{\perp}\subset\mathfrak{g}^{*}$ be the annihilator of $\mathfrak{u}_{P}$. Then $T^{*}(G/U_{P})$ can be identified with $G\times_{U_{P}}\mathfrak{u}_{P}^{\perp}$. Let $E$ be the homogeneous vector bundle $G\times_{B}\mathfrak{u}_{P}^{\perp}$ over $G/B$.
Theorem 1.6.
Let $G$ be a connected semi-simple group, and let $U_{P}$ be the unipotent radical of a parabolic subgroup $P$ of $G$.
  1. (1.6.1)
    Then $S(G/U_{P})$ is finitely generated if and only if $\mathbb{P}E^{*}$ is a Mori dream space.
  2. (1.6.2)
    The variety $\overline{T^{*}(G/U_{P})}$ is symplectic if and only if $\mathbb{P}E^{*}$ is of Fano type.
For the definitions of Mori dream spaces and varieties of Fano type, we refer the reader to Sections 2.1 and 2.2. Based on Theorem 1.6, we propose the following conjecture, which is supported by the evidence provided in Proposition 5.5 (see also Corollary 5.2).
Conjecture 1.7.
The projective bundle $\mathbb{P}E^{*}$ is of Fano type.
Acknowledgments
We would like to thank B. Jia for useful discussions and I. Losev for drawing our attention to the works of T. Gannon. We are grateful to A. Beauville, T. Gannon and S. Takagi for providing valuable feedback on the preliminary version of this paper. J. Liu is supported by the National Key Research and Development Program of China (No. 2021YFA1002300) and the Youth Innovation Promotion Association CAS. Both authors are supported by the CAS Project for Young Scientists in Basic Research (No. YSBR-033) and the NSFC grant (No. 12288201).
2. Preliminaries
All varieties are assumed to be integral schemes of finite type over $k$, and all groups are assumed to be affine algebraic groups over $k$.
2.1. Cox rings and Mori dream spaces
Let $X$ be a normal variety with a finitely generated free divisor class group $\operatorname{Cl}(X)$. Let $D_{1},\dots,D_{\rho}$ be Weil divisors whose classes form a $\mathbb{Z}$-basis of $\operatorname{Cl}(X)$. The Cox ring of $X$ is defined as the following $\operatorname{Cl}(X)$-graded ring:
$\mathcal{R}(X)\coloneqq\bigoplus_{(m_{1},\dots,m_{\rho})\in\mathbb{Z}^{\rho}}H^{0}(X,\scr{O}_{X}(m_{1}D_{1}+\dots+m_{\rho}D_{\rho})).$
This definition can be extended to normal varieties $X$ with a finitely generated divisor class group $\operatorname{Cl}(X)$ that may have torsion, provided that $H^{0}(X,\scr{O}_{X}^{\times})=k^{\times}$. For further details on this construction, we refer the reader to [ADHL15, § 1.4.2].
Lemma 2.1.
Let $X$ be a normal variety such that $H^{0}(X,\scr{O}_{X}^{\times})=k^{\times}$ and $\operatorname{Cl}(X)$ is finitely generated. Let $\mathbb{T}\coloneqq\operatorname{Spec}(k[\operatorname{Cl}(X)])$ be the quasitorus acting naturally on $\mathcal{R}(X)$. Then $k[X]=\mathcal{R}(X)^{\mathbb{T}}$. In particular, if $\operatorname{Cl}(X)=0$, then $k[X]=\mathcal{R}(X)$.
Proof.
This follows from [ADHL15, Example 1.2.3.4]. ∎
Theorem 2.2 ([EKW04, Corollary 1.2]).
Let $X$ be a normal variety such that $\operatorname{Cl}(X)$ is torsion-free and finitely generated. Then $\mathcal{R}(X)$ is a unique factorization domain (UFD).
The following result is a special case of [ADHL15, Theorem 4.4.1.2], and it will play a central role in the proof of our main results.
Theorem 2.3.
Let $X$ be a smooth variety with finitely generated $\operatorname{Cl}(X)$ and $H^{0}(X,\scr{O}_{X}^{\times})=k^{\times}$. Assume that there exists a free algebraic torus action $\mathbb{T}\times X\rightarrow X$ admitting a smooth geometric quotient $X/\mathbb{T}$. Then $H^{0}(X/\mathbb{T},\scr{O}_{X/\mathbb{T}}^{\times})=k^{\times}$, $\operatorname{Cl}(X/\mathbb{T})$ is finitely generated, and there is a $\operatorname{Cl}(X)$-graded isomorphism $\mathcal{R}(X/\mathbb{T})\rightarrow\mathcal{R}(X)$.
Definition 2.4 ([HK00]).
A Mori dream space is an irreducible, normal projective variety $X$ with finitely generated $\operatorname{Cl}(X)$ and finitely generated Cox ring $\mathcal{R}(X)$.
2.2. Varieties of Fano type
A log pair $(X,\Delta)$ consists of a normal variety $X$ and an effective $\mathbb{Q}$-divisor $\Delta$ on $X$ such that $K_{X}+\Delta$ is $\mathbb{Q}$-Cartier. For the definitions of terminal, canonical, and klt singularities of a log pair, we refer to [KM98, § 2.3]. In particular, a normal variety $X$ is said to have terminal, canonical, or klt singularities if the log pair $(X,0)$ has the corresponding property. Moreover, a normal variety $X$ has klt type singularities if there exists an effective $\mathbb{Q}$-divisor $\Delta$ on $X$ such that $K_{X}+\Delta$ is $\mathbb{Q}$-Cartier and $(X,\Delta)$ is klt. The following result follows directly from the definitions and [KM98, Theorem 5.22 and Corollary 5.24].
Lemma 2.5.
Let $X$ be a normal variety such that $K_{X}$ is Cartier. Then the following statements are equivalent:
  1. (2.5.1)
    $X$ has rational singularities.
  2. (2.5.2)
    $X$ has canonical singularities.
  3. (2.5.3)
    $X$ has klt singularities.
  4. (2.5.4)
    $X$ has klt type singularities.
Recall that a projective manifold is called Fano (resp. weak Fano) if $-K_{X}$ is ample (resp. big and nef).
Definition 2.6.
A normal projective variety $X$ is said to be of Fano type if there exists an effective $\mathbb{Q}$-divisor $\Delta$ such that $-(K_{X}+\Delta)$ is an ample $\mathbb{Q}$-Cartier divisor and the log pair $(X,\Delta)$ has klt singularities.
It is known that a weak Fano manifold is of Fano type. Moreover, varieties of Fano type can be characterized in terms of the singularities of their Cox rings.
Theorem 2.7 ([GOST15, Theorem 1.1 and Corollary 1.3]).
Let $X$ and $Y$ be $\mathbb{Q}$-factorial normal projective varieties with finitely generated divisor class groups.
  1. (2.7.1)
    The variety $X$ is of Fano type if and only if $\mathcal{R}(X)$ is finitely generated and $\operatorname{Spec}(\mathcal{R}(X))$ has klt singularities.
  2. (2.7.2)
    If there exists a surjective morphism $f\colon X\rightarrow Y$ and $X$ is of Fano type, then $Y$ is also of Fano type.
2.3. Invariant-theoretic quotients
Let $A$ be a ring, and let $G$ be a group acting on $A$. Denote by $A^{G}$ the invariant subring of $A$. It is well-known that for an affine algebraic variety $X=\operatorname{Spec}A$ over $k$ and $G$ a reductive group, the invariant ring $A^{G}$ is finitely generated, and the (affine invariant-theoretic) quotient $X/\!\!/G$ is defined to be the affine variety $\operatorname{Spec}A^{G}$. The following result relates the properties of $X$ to its quotient — see [KM98, Lemma 5.16], [Zhu24, Corollary 1.2], and [BGLM24, Theorem 1] for more details.
Theorem 2.8.
Let $X=\operatorname{Spec}A$ be a normal affine algebraic variety over $k$, and let $G$ be a group acting on $X$. Set $Y\coloneqq X/\!\!/G$.
  1. (2.8.1)
    If $G$ is finite and $X$ is $\mathbb{Q}$-factorial, then $Y$ is $\mathbb{Q}$-factorial.
  2. (2.8.2)
    If $G$ is reductive and $X$ has klt singularities, then $Y$ has klt type singularities. In particular, if $K_{Y}$ is $\mathbb{Q}$-Cartier, then $Y$ has klt singularities.
In general, the invariant ring under a non-reductive group action is not necessarily finitely generated. However, we have the following interesting result from [Gro97, Theorem 16.2].
Proposition 2.9.
Let $A$ be a commutative $k$-algebra on which a reductive group $G$ acts rationally. Let $U\subset G$ be a maximal unipotent subgroup. Then $A$ is finitely generated if and only if $A^{U}$ is finitely generated.
Theorem 2.10.
Let $G$ be a connected reductive group, and let $U_{P}$ be the unipotent radical of a parabolic subgroup $P$ of $G$. Let $A$ be a commutative $k$-algebra on which $G$ acts rationally. Then $A$ is finitely generated (resp. integrally closed, or has rational singularities) if and only if $A^{U_{P}}$ has the same property.
Proof.
Let $B$ be the Borel subgroup contained in $P$. Let $p\colon G^{\prime}\rightarrow G$ be an isogeny such that $G^{\prime}$ is of simply connected type, i.e., $G^{\prime}=G_{\text{ss}}\times C$, where $G_{\text{ss}}$ is simply connected semi-simple and $C$ is a torus. Then there exists a parabolic subgroup $P_{\text{ss}}$ of $G_{\text{ss}}$ such that $p(P^{\prime})=P$ and $p(B^{\prime})=B$, where $P^{\prime}=P_{\text{ss}}\times C$, $B^{\prime}=B_{\text{ss}}\times C$, and $B_{\text{ss}}$ is a Borel subgroup of $G_{\text{ss}}$ contained in $P_{\text{ss}}$. Then $p(U_{B_{\text{ss}}})=U_{B}$ and $p(U_{P_{\text{ss}}})=U_{P}$, so after replacing $G$ by $G_{\text{ss}}$, we may assume that $G$ itself is simply connected semi-simple. Applying Proposition 2.9 and [Pop86, Theorem 6 and its corollary] to the action of $G$ on $A$ shows that $A$ has the required properties in the theorem if and only if $A^{U_{B}}$ does.
Let $L$ be the Levi subgroup of $P$ with unipotent radical $U_{L}$. Then $A^{U_{B}}=(A^{U_{P}})^{U_{L}}$. Applying the previous argument to the action of $L$ on $A^{U_{P}}$ yields that $A^{U_{P}}$ has the properties in the statement if and only if $A^{U_{B}}$ does, and hence if and only if $A$ does, as desired. ∎
Lemma 2.11 ([Gro97, Lemma 20.1]).
Let $A$ be a commutative $k$-algebra, and let $G$ be a connected group acting on $A$ rationally. If $A$ is a UFD such that its group of units is $k^{\times}$ and $G$ has no non-trivial character, then $A^{G}$ is again a UFD.
Remark 2.12.
Recall from [Har77, Proposition II.6.2] that a Noetherian domain $A$ is a UFD if and only if $\operatorname{Spec}(A)$ is normal with a trivial divisor class group.
3. The affine closure of cotangent bundles
3.1. Preliminary observations
As is well-known, the $k$-algebra $k[X]$ of regular functions on a smooth variety $X$ is not necessarily finitely generated. In the case where $X$ is quasi-affine, we have the following:
Proposition 3.1.
Let $X$ be a smooth quasi-affine variety.
  1. (3.1.1)
    If $k[X]$ is finitely generated, then the natural morphism $X\rightarrow\overline{X}$ is an open embedding with a small boundary, i.e., $\operatorname{codim}(\overline{X}\setminus X)\geq 2$.
  2. (3.1.2)
    The ring $k[X]$ is finitely generated if and only if $X$ embeds into an affine variety with a small boundary, i.e., $X=Z\setminus F$, where $Z$ is an affine variety and $F\subset Z$ is a closed subvariety of codimension at least two.
  3. (3.1.3)
    If $X$ is an open subset of a factorial affine variety, then $k[X]$ is finitely generated.
Proof.
Statement (3.1.1) follows from [Gro97, Theorem 4.2].
For (3.1.2), assume $X=Z\setminus F$ for an affine variety $Z$ with $\operatorname{codim}F\geq 2$. Then $X$ embeds into the normalization $\tilde{Z}$ with a small boundary, implying $k[X]=k[\tilde{Z}]$ is finitely generated.
For (3.1.3), assume $X\subset Z$ is open with $Z$ a factorial affine variety. Since $Z\setminus D$ is affine for any irreducible divisor $D\subset Z$, we apply (3.1.2) by removing divisors in $Z\setminus X$. ∎
Recall that for a smooth variety $X$, its algebra of symmetric tensors $S(X)$ is naturally isomorphic to $k[T^{*}X]$.
Proposition 3.2.
Let $X$ be a smooth quasi-affine variety. Then $\overline{T^{*}X}$ is a symplectic variety if and only if $S(X)$ is finitely generated and $\overline{T^{*}X}$ has rational (equivalently canonical or klt or klt type) singularities.
Proof.
If $S(X)$ is finitely generated, the complement of the open embedding $\varphi_{X}\colon T^{*}X\to\overline{T^{*}X}$ has codimension $\geq 2$ by Proposition 3.1. Thus, the symplectic structure on $T^{*}X$ extends to $(\overline{T^{*}X})_{\text{reg}}$, implying the canonical divisor of $\overline{T^{*}X}$ is trivial. If $\overline{T^{*}X}$ has rational (or canonical/klt/klt type) singularities, it is symplectic by Lemma 2.5 and [Nam01, Theorem 6]. ∎
Proposition 3.3.
Let $X$ be an open subset of a smooth affine variety. Then the affine closures $\overline{X}$ and $\overline{T^{*}X}$ are both smooth, and $\overline{T^{*}X}\cong T^{*}\overline{X}$.
Proof.
Write $X\coloneqq Z\setminus F$ for a smooth affine variety $Z$ and closed subset $F\subset Z$. If $F$ is an irreducible divisor, $Z\setminus F$ is affine, so $\overline{X}=X$ and $\overline{T^{*}X}=T^{*}X$. If $\operatorname{codim}F\geq 2$, then $\overline{X}=Z$ and $\overline{T^{*}X}=T^{*}Z$. The result follows by induction. ∎
Example 3.4.
Let $M$ be a smooth affine variety on which a finite group $\Gamma$ acts freely on an open subset $M_{0}$ with $\operatorname{codim}(M\setminus M_{0})\geq 2$. Let $Z=M/\Gamma$ be the quotient, and let $X\subset Z$ be its smooth locus. Then $X$ contains $M_{0}/\Gamma$ with a boundary of codimension $\geq 2$. Hence, $k[T^{*}X]\cong k[T^{*}(M_{0}/\Gamma)]$, and since $T^{*}(M_{0}/\Gamma)$ is an open subset of $(T^{*}M)/\Gamma$ with codimension $\geq 2$ boundary, we have $\overline{T^{*}X}\cong(T^{*}M)/\Gamma$.
The following results show that we can freely pass to finite étale covers:
Lemma 3.5.
Let $f\colon X\rightarrow Y$ be a finite morphism between smooth quasi-affine varieties. If $k[Y]$ is finitely generated, then the induced morphism $\overline{X}\rightarrow\overline{Y}$ is finite. In particular, $k[X]$ is finitely generated.
Proof.
Let $\iota\colon Y\rightarrow\overline{Y}$ be the natural open embedding. Since $f$ is proper, $f_{*}\scr{O}_{X}$ is coherent. As $\operatorname{codim}(\overline{Y}\setminus Y)\geq 2$ by Proposition 3.1, the sheaf $\iota_{*}f_{*}\scr{O}_{X}$ is coherent. Therefore, $k[X]=H^{0}(Y,f_{*}\scr{O}_{X})=H^{0}(\overline{Y},\iota_{*}f_{*}\scr{O}_{X})$ is a finite $k[Y]$-module. ∎
Proposition 3.6.
Let $f\colon X\rightarrow Y$ be an étale Galois morphism between smooth quasi-affine varieties. Then:
  1. (3.6.1)
    $S(X)$ is finitely generated if and only if $S(Y)$ is finitely generated.
  2. (3.6.2)
    $\overline{T^{*}X}$ is symplectic if and only if $\overline{T^{*}Y}$ is symplectic.
  3. (3.6.3)
    If $S(X)$ is finitely generated and $\overline{T^{*}X}$ is $\mathbb{Q}$-factorial, then $\overline{T^{*}Y}$ is $\mathbb{Q}$-factorial.
Proof.
The induced morphism $T^{*}X\rightarrow T^{*}Y$ is an étale Galois cover, so (3.6.1) follows from Lemma 3.5. If $S(X)$ and $S(Y)$ are finitely generated, then $\overline{T^{*}X}\rightarrow\overline{T^{*}Y}$ is a quasi-étale finite Galois cover. Statements (3.6.2) and (3.6.3) follow from [KM98, Proposition 5.20], Proposition 3.2 and Theorem 2.8. ∎
3.2. Examples of non-finite generation
We now give examples of smooth varieties whose algebra of symmetric tensors is not finitely generated.
Example 3.7.
Let $X$ be a smooth variety such that $k[X]$ is not finitely generated over $k$. Let $\sigma\colon X\rightarrow T^{*}X$ be the zero section of $\pi\colon T^{*}X\rightarrow X$. Then the following composition of morphisms of $k$-algebras
$k[X]\stackrel{{\scriptstyle\pi^{*}}}{{\longrightarrow}}k[T^{*}X]\stackrel{{\scriptstyle\sigma^{*}}}{{\longrightarrow}}k[X]$
coincides with the identity. This implies that $k[T^{*}X]$ is not finitely generated.
Example 3.8.
Let $Y$ be a smooth projective variety with a nef and big line bundle $L$ on $Y$ such that the ring $R(Y,L)$ of sections is not finitely generated. Let $E$ be the rank two vector bundle $L^{-1}\oplus L$ over $Y$. Denote by $X$ the projective bundle $\pi\colon\mathbb{P}(E)\rightarrow Y$ and by $\sigma\colon Y\rightarrow\mathbb{P}(E)$ the section corresponding to the natural quotient $E\rightarrow L$. Let $Z\coloneqq\sigma(Y)$. Then we have $T_{X}|_{Z}\cong T_{Z}\oplus T_{X/Y}|_{Z}$, which yields
$\sigma^{*}T_{X}\cong\sigma^{*}T_{Z}\oplus\sigma^{*}T_{X/Y}\cong T_{Y}\oplus\sigma^{*}\scr{O}_{\mathbb{P}(E)}(2)\cong T_{Y}\oplus L^{\otimes 2}.$
Consider the following composition of morphisms of $k$-algebras
$q\colon k[T^{*}X]\cong S(X)\longrightarrow\bigoplus_{k\geq 0}H^{0}(Y,{\sf S}^{k}L^{\otimes 2})=R(Y,L^{\otimes 2}).$
For any integer $k\geq 0$, the natural morphism
${\sf S}^{2k}E=\pi_{*}\scr{O}_{\mathbb{P}(E)}(2k)=\sigma^{*}\pi^{*}\pi_{*}\scr{O}_{\mathbb{P}(E)}(2k)\rightarrow\sigma^{*}\scr{O}_{\mathbb{P}(E)}(2k)\cong L^{\otimes 2k}$
coincides with the natural projection ${\sf S}^{2k}E\rightarrow{\sf S}^{2k}L=L^{\otimes 2k}$. Thus, the natural inclusion
$\bigoplus_{k\geq 0}H^{0}(Y,{\sf S}^{2k}E)=\bigoplus_{k\geq 0}H^{0}(X,T_{X/Y}^{\otimes k})\longrightarrow k[T^{*}X]$
shows that $q$ is surjective, and hence $k[T^{*}X]$ is not finitely generated.
4. Braverman–Kazhdan and horospherical spaces
4.1. Recollections on homogeneous spaces
We recall the basic facts about algebraic groups and homogeneous spaces.
Lemma 4.1.
Let $G$ be a connected algebraic group, and let $H$ be a closed subgroup of $G$. Denote by $\mathbb{X}(G)$ (resp. $\mathbb{X}(H)$) the character group of $G$ (resp. $H$).
  1. (4.1.1)
    If $\mathbb{X}(H)=0$, then $H$ is observable, i.e., $G/H$ is quasi-affine.
  2. (4.1.2)
    If $G$ is simply connected semi-simple, then $k[G]$ is a UFD.
  3. (4.1.3)
    If $\mathbb{X}(G)=0$, then $H^{0}(G/H,\mathcal{O}_{G/H}^{\times})=k^{\times}$.
Proof.
Statements (4.1.1) and (4.1.2) follow from [Gro97, Corollary 1.5 and Theorem 20.6], respectively. For statement (4.1.3), note that $H^{0}(G/H,\scr{O}_{G/H})$ is a subalgebra of $H^{0}(G,\scr{O}_{G})$, so it suffices to prove $H^{0}(G,\scr{O}_{G}^{\times})=k^{\times}$. Let $f\colon G\rightarrow k^{\times}$ be an invertible function. Up to scaling, we may assume $f(e)=1$. Then Rosenlicht’s unit theorem ([KKV89, Proposition 1.2]) implies that $f$ is a character of $G$ and hence a constant, as desired. ∎
Lemma 4.2.
Let $G$ be a connected algebraic group, and let $H$ be a closed subgroup.
  1. (4.2.1)
    The group $\operatorname{Cl}(G)$ is finite, and if $[G,G]$ is simply connected, then $\operatorname{Cl}(G)=0$.
  2. (4.2.2)
    The group $\operatorname{Cl}(G/H)$ is finitely generated, and there exists an exact sequence
    $\mathbb{X}(G)\longrightarrow\mathbb{X}(H)\longrightarrow\operatorname{Cl}(G/H)\longrightarrow\operatorname{Cl}(G).$
Proof.
See, for example, [ADHL15, Theorem 4.2.2.1, Theorem 4.5.1.2, and Corollary 4.5.1.4] and [Pop23, Corollary]. ∎
4.2. Proof of Theorem 1.2
In the sequel of this section, we will denote by $X$ a horospherical homogeneous space $G/H$. Then $T^{*}X$ can be canonically identified with the homogeneous bundle $G\times_{H}\mathfrak{h}^{\perp}$, where $\mathfrak{h}^{\perp}\subset\mathfrak{g}^{*}$ is the annihilator of $\mathfrak{h}$. As $P=N_{G}(H)$, we have a $G$-equivariant map
$\phi\colon T^{*}X\cong G\times_{H}\mathfrak{h}^{\perp}\longrightarrow G\times_{P}\mathfrak{h}^{\perp}\eqqcolon E.$
Then $E$ is a homogeneous bundle over the homogeneous space $G/P$.
Lemma 4.3.
The Cox ring $\mathcal{R}(\mathbb{P}E^{*})$ is finitely generated, and $\operatorname{Spec}(\mathcal{R}(\mathbb{P}E^{*}))$ is a factorial variety with canonical Gorenstein singularities.
Proof.
By [FL24, Lemma 5.3], the projective manifold $\mathbb{P}E^{*}$ is weak Fano and hence of Fano type. It follows from Theorem 2.2 and Theorem 2.7 that $\mathcal{R}(\mathbb{P}E^{*})$ is a finitely generated UFD and $\operatorname{Spec}(\mathcal{R}(\mathbb{P}E^{*}))$ has klt singularities. In particular, the canonical bundle of $\operatorname{Spec}(\mathcal{R}(\mathbb{P}E^{*}))$ is Cartier, hence $\operatorname{Spec}(\mathcal{R}(\mathbb{P}E^{*}))$ is Gorenstein with canonical singularities by Lemma 2.5. ∎
Lemma 4.4.
Let $X=G/H$ be a horospherical homogeneous space, and denote by $\bar{\pi}\colon\mathbb{P}T_{X}\to X$ the natural projection. The affinization morphism $\varphi_{X}\colon T^{*}X\rightarrow\overline{T^{*}X}$ is birational.
Proof.
Let $Z$ be a smooth $G$-equivariant completion of $X$. Then the natural restriction $S(Z)\rightarrow S(X)$ induces a commutative diagram
[…]
Since $Z$ is a smooth projective horospherical variety, the tangent bundle $T_{Z}$ is big by [Liu23, Theorem 1.2]. Thus, the map $\varphi_{Z}$ is birational by [FL24, Theorem 1.6]. As $T^{*}X\rightarrow T^{*}Z$ is an open embedding and $\overline{T^{*}X}\rightarrow\overline{T^{*}Z}$ is dominant, it follows that $\varphi_{X}$ is also birational. ∎
Now we are in the position to prove Theorem 1.2, which is a direct consequence of the following slightly more general result.
Theorem 4.5.
Let $G$ be a connected semi-simple group, and let $X\coloneqq G/H$ be a horospherical homogeneous space.
  1. (4.5.1)
    The $k$-algebra $S(X)$ is finitely generated, and the affinization $\overline{T^{*}X}$ has klt type singularities.
  2. (4.5.2)
    The affinization $\overline{T^{*}X}$ is factorial if $\operatorname{Cl}(X)=0$ and $\mathbb{Q}$-factorial if $\operatorname{Cl}(X)$ is finite.
  3. (4.5.3)
    If $X$ is quasi-affine, then $\overline{T^{*}X}$ is a symplectic variety.
Proof.
Since $G$ is semi-simple, $\mathbb{X}(G)=0$, hence $H^{0}(X,\scr{O}_{X}^{\times})=k^{\times}$ by Lemma 4.1, which implies $H^{0}(\mathbb{P}T_{X},\mathcal{O}_{\mathbb{P}T_{X}}^{\times})=k^{\times}$.
Thanks to Theorem 2.3 and Lemma 4.4, the geometric quotient $\mathbb{P}T_{X}\rightarrow\mathbb{P}E^{*}$ induces a graded isomorphism $\mathcal{R}(\mathbb{P}T_{X})\cong\mathcal{R}(\mathbb{P}E^{*})$. Moreover, since the cokernel of the homomorphism $\mathbb{Z}\scr{O}_{\mathbb{P}T_{X}}(1)\rightarrow\operatorname{Cl}(\mathbb{P}T_{X})$ is naturally isomorphic to $\operatorname{Cl}(X)$ by Lemma 4.4, the characteristic quasitorus $\mathbb{T}\coloneqq\operatorname{Spec}(k[\operatorname{Cl}(X)])$ acts naturally on $\mathcal{R}(\mathbb{P}T_{X})$ such that
$\mathcal{R}(\mathbb{P}T_{X})^{\mathbb{T}}=\bigoplus_{p\geq 0}H^{0}(\mathbb{P}T_{X},\scr{O}_{\mathbb{P}T_{X}}(p))\cong\bigoplus_{p\geq 0}H^{0}(X,{\sf S}^{p}T_{X})=S(X).$
Thus, $S(X)$ is finitely generated, and $\overline{T^{*}X}$ has klt type singularities by Theorem 2.8 and Lemma 4.3. If $\operatorname{Cl}(X)=0$, then $S(X)=\mathcal{R}(\mathbb{P}E^{*})$ is a UFD, hence $\overline{T^{*}X}$ is factorial. If $\operatorname{Cl}(X)$ is finite, so is $\mathbb{T}$, and thus $\overline{T^{*}X}$ is $\mathbb{Q}$-factorial by Theorem 2.8.
Finally, it follows from Proposition 3.2 that $\overline{T^{*}X}$ has symplectic singularities if $X$ is quasi-affine. ∎
Remark 4.6.
If $X=G/H$ is not quasi-affine, the preceding argument shows that if $\varphi_{X}(T^{*}X)$ contains an open subset $V$ of $\overline{T^{*}X}$ such that $\operatorname{codim}(\overline{T^{*}X}\setminus V)\geq 2$, then $\overline{T^{*}X}$ is a symplectic variety.
Using the same proof, we can provide an answer to Question 1.4 for projective rational homogeneous spaces as follows.
Proposition 4.7.
Let $Z\cong G/P\subset\mathbb{P}^{n}$ be a $G$-equivariant embedding of a rational homogeneous space, and let $\hat{Z}\subset\mathbb{A}^{n+1}$ be the affine cone. Consider the quasi-affine variety $X=\hat{Z}\setminus\{0\}$. Then $\overline{T^{*}X}$ is a symplectic variety.
Proof.
Let $\hat{G}\subset\operatorname{GL}_{n+1}$ be the pre-image of $G\subset\operatorname{PGL}_{n+1}$. Then we have an exact sequence $1\to k^{\times}\to\hat{G}\to G\to 1$. The variety $X$ is homogeneous under the natural action of $\hat{G}$ with stabilizers isomorphic to $P$. It follows that $X$ is horospherical, and we can apply Theorem 1.2 to conclude. ∎
Proof of Corollary 1.3.
Let $p\colon G^{\prime}\rightarrow G$ be an isogeny such that $G^{\prime}=G_{\text{ss}}\times C$, where $G_{\text{ss}}$ is simply connected semi-simple and $C$ is a torus. Then there exists a parabolic subgroup $P_{\text{ss}}$ of $G_{\text{ss}}$ such that $p(P^{\prime})=P$, where $P^{\prime}=P_{\text{ss}}\times C$, and the induced morphism
$G^{\prime}/[P^{\prime},P^{\prime}]=G_{\text{ss}}/[P_{\text{ss}},P_{\text{ss}}]\times C\longrightarrow G/[P,P]$
is an étale Galois cover. Notice that $T^{*}C=C\times\mathbb{A}^{r}$ is a smooth affine variety, where $r=\dim(C)$, and the algebra $S(G_{\text{ss}}/[P_{\text{ss}},P_{\text{ss}}])$ is finitely generated by Theorem 4.5. Thus, the algebra
$S(G^{\prime}/[P^{\prime},P^{\prime}])\cong S(G_{\text{ss}}/[P_{\text{ss}},P_{\text{ss}}])\otimes S(C)$
is finitely generated (cf. [BL24, Lemma 4.2]), and we have
$\overline{T^{*}(G^{\prime}/[P^{\prime},P^{\prime}])}\cong\overline{T^{*}(G_{\text{ss}}/[P_{\text{ss}},P_{\text{ss}}])}\times T^{*}C.$
Since $\mathbb{X}([P_{\text{ss}},P_{\text{ss}}])=0$, the variety $G_{\text{ss}}/[P_{\text{ss}},P_{\text{ss}}]$ is quasi-affine, and Lemma 4.2 yields $\operatorname{Cl}(G_{\text{ss}}/[P_{\text{ss}},P_{\text{ss}}])=0$. Thus Theorem 1.2 implies that $\overline{T^{*}(G_{\text{ss}}/[P_{\text{ss}},P_{\text{ss}}])}$ is a factorial symplectic variety, and hence $\overline{T^{*}(G^{\prime}/[P^{\prime},P^{\prime}])}$ is factorial and symplectic.
Finally, it follows from Proposition 3.6 that the algebra $S(G/[P,P])$ is finitely generated, and $\overline{T^{*}X}$ is a $\mathbb{Q}$-factorial symplectic variety. Moreover, if $[G,G]$ is simply connected, then $\operatorname{Cl}(G)=\operatorname{Cl}(G/[P,P])=0$ by Lemma 4.2 as $\mathbb{X}([P,P])=0$, which implies
$\operatorname{Cl}(\overline{T^{*}(G/[P,P])})=\operatorname{Cl}(T^{*}(G/[P,P]))=\operatorname{Cl}(G/[P,P])=0.$
Hence $\overline{T^{*}(G/[P,P])}$ is factorial by Remark 2.12. ∎
4.3. Associated varieties
Let $G$ be a connected simple group, and let $P\subset G$ be a maximal parabolic subgroup containing a Borel subgroup $B$. Let $\mathfrak{h}\subset\mathfrak{b}$ be the Cartan subalgebra. Let $\mathfrak{g}=\mathfrak{n}^{+}\oplus\mathfrak{h}\oplus\mathfrak{n}^{-}$ be the usual triangular decomposition relative to the choice of $\mathfrak{h}$ and $\mathfrak{b}=\mathfrak{h}\oplus\mathfrak{n}^{+}$. Write $\mathfrak{p}=\mathfrak{l}\oplus\mathfrak{u}^{+}$, where $\mathfrak{u}^{+}$ is the nilpotent radical of $\mathfrak{p}$ and $\mathfrak{l}$ is the Levi part. Thus, $\mathfrak{g}=\mathfrak{u}^{+}\oplus\mathfrak{l}\oplus\mathfrak{u}^{-}$. We will consider certain cases where $\mathfrak{u}^{+}$ is abelian, summarized in the following table (see [LSS88, Table 3.1] and [LS89, Lemmas, p. 32 and p. 37]).
Table 1.
Type of $\mathfrak{g}$ Dynkin diagram parabolics $\mathfrak{p}$ with abelian radical nilpotent orbit $\mathcal{O}$ $\dim(\mathcal{O})$ $\dim(X)$
[…]
$E_{6}$dynkin$\dynkin[labels={1,...,6}]E6$ $\mathfrak{p}_{1}$, $\mathfrak{p}_{6}$ $\mathcal{O}_{\min}$ $22$ $11$
$E_{7}$dynkin$\dynkin[labels={1,...,7}]E7$ $\mathfrak{p}_{7}$ $\mathcal{O}_{\min}$ $34$ $17$
The extra notation in Table LABEL:table.Abelian is as follows. In each case, $\mathfrak{p}_{i}$ is the maximal parabolic subalgebra of $\mathfrak{g}$ associated to the simple root $\alpha_{i}$ (in the order of Bourbaki), and $\mathcal{O}_{\min}$ denotes the minimal non-zero nilpotent orbit in $\mathfrak{g}$. In Cases $A_{n}$, $C_{n}$, and $D_{n}$, the variety $\mathcal{O}_{r}$ is a nilpotent orbit in $\mathfrak{g}$ defined as follows (see [LS89, II, § 5 and § 6]):
$\mathcal{O}_{r}\coloneqq\{z\in\mathfrak{g}\mid\operatorname{rank}(z)=\widetilde{r}\text{ and }z^{2}=0\},$
where $\widetilde{r}=r$ in Cases $A_{n}$ and $C_{n}$, but $\widetilde{r}=2r$ in Cases $D_{n}$. Once both $\mathfrak{p}$ and $\mathcal{O}$ are fixed, we define $X$ to be $\mathcal{O}\cap\mathfrak{u}^{+}$. According to [LSS88, Remarks 3.2] and [LS89, Lemmas, p. 32 and p. 37], the affine closure of $X$ (resp. $\mathcal{O}$) is exactly the closure of $X$ (resp. $\mathcal{O}$) in $\mathfrak{u}^{+}$ (resp. $\mathfrak{g}$).
Now consider the $k^{\times}$-action on $\mathfrak{g}$ with weight 00 (resp. $1$ and $2$) on $\mathfrak{u}^{+}$ (resp. $\mathfrak{l}$ and $\mathfrak{u}^{-}$), namely $\phi:k^{\times}\times\mathfrak{g}\to\mathfrak{g}$ such that
$\phi_{\lambda}(u^{+}+l+u^{-})=u^{+}+\lambda l+\lambda^{2}u^{-}.$
Lemma 4.8.
For any $v,x,y\in\mathfrak{g}$, we have:
  1. (4.8.1)
    $[\phi_{\lambda}(x),\phi_{\lambda}(y)]=\lambda\phi_{\lambda}([x,y])$.
  2. (4.8.2)
    $\kappa(\phi_{\lambda}(v),[\phi_{\lambda}(x),\phi_{\lambda}(y)])=\lambda^{3}\kappa(v,[x,y])$, where $\kappa$ is the Killing form.
  3. (4.8.3)
    Any nilpotent orbit in $\mathfrak{g}$ is stable under this $k^{\times}$-action.
Proof.
Write $x=x^{+}+x_{0}+x^{-}$ (resp. $y=y^{+}+y_{0}+y^{-}$) in the decomposition of $\mathfrak{g}=\mathfrak{u}^{+}\oplus\mathfrak{l}\oplus\mathfrak{u}^{-}$. Recall that $\mathfrak{u}^{+}$ and $\mathfrak{u}^{-}$ are both abelian, hence we have
$[\phi_{\lambda}(x),\phi_{\lambda}(y)]$ $=[x^{+}+\lambda x_{0}+\lambda^{2}x^{-},y^{+}+\lambda y_{0}+\lambda^{2}y^{-}]$
$=\lambda([x^{+},y_{0}]+[x_{0},y^{+}])+\lambda^{2}([x^{+},y^{-}]+[x_{0},y_{0}]+[x^{-},y^{+}])$
$\quad+\lambda^{3}([x_{0},y^{-}]+[x^{-},y_{0}])$
$=\lambda\phi_{\lambda}([x,y]).$
Now consider the map $\psi_{\lambda}=\frac{1}{\lambda}\phi_{\lambda}$. By (4.8.1), $\psi_{\lambda}$ preserves the Lie bracket. This implies that $\psi_{\lambda}$ preserves the Killing form, hence we have
$\kappa(\phi_{\lambda}(v),[\phi_{\lambda}(x),\phi_{\lambda}(y)])$ $=\kappa(\phi_{\lambda}(v),\lambda\phi_{\lambda}([x,y]))$
$=\kappa(\lambda\psi_{\lambda}(v),\lambda^{2}\psi_{\lambda}([x,y]))$
$=\lambda^{3}\kappa(v,[x,y]).$
Since $\psi_{\lambda}$ preserves the Lie bracket, its action on $\mathfrak{g}$ is via the adjoint action, hence it preserves any adjoint orbits in $\mathfrak{g}$. As nilpotent orbits are stable under the dilation action of scalars, it follows that $\phi_{\lambda}$ preserves nilpotent orbits. ∎
Theorem 4.9.
Let $\mathfrak{g}$ be a simple Lie algebra with $\mathfrak{p}$, $\mathcal{O}$, and $X$ as in Table LABEL:table.Abelian. Then $\overline{T^{*}X}$ is isomorphic to $\overline{\mathcal{O}}\subset\mathfrak{g}$.
Proof.
For an element $v\in\mathcal{O}$, the tangent space $T_{v}\mathcal{O}$ is naturally identified with $[v,\mathfrak{g}]$. By definition, the symplectic form on $\mathcal{O}$ is given by $\omega_{v}([v,x],[v,y])=\kappa(v,[x,y])$.
We now show that $\phi_{\lambda}^{*}\omega=\lambda\omega$ for all $\lambda\in k^{\times}$. Take any $v\in\mathcal{O}$ and $x,y\in\mathfrak{g}$; then
$(\phi_{\lambda}^{*}\omega)_{v}([v,x],[v,y])$ $=\omega_{\phi_{\lambda}(v)}(d\phi_{\lambda}([v,x]),d\phi_{\lambda}([v,y]))$
$=\omega_{\phi_{\lambda}(v)}\left([\phi_{\lambda}(v),\frac{1}{\lambda}\phi_{\lambda}(x)],[\phi_{\lambda}(v),\frac{1}{\lambda}\phi_{\lambda}(y)]\right)$
$=\kappa\left(\phi_{\lambda}(v),\frac{1}{\lambda^{2}}[\phi_{\lambda}(x),\phi_{\lambda}(y)]\right)$
$=\frac{1}{\lambda^{2}}\kappa(\phi_{\lambda}(v),[\phi_{\lambda}(x),\phi_{\lambda}(y)])$
$=\lambda\kappa(v,[x,y])=\lambda\omega_{v}([v,x],[v,y]).$
Notice that $\dim X=\frac{1}{2}\dim\mathcal{O}$, and the $k^{\times}$-fixed locus is $\mathcal{O}\cap\mathfrak{u}^{+}=X$. It follows that $X$ is a Lagrangian submanifold of $\mathcal{O}$. Consider the map $\pi_{+}\colon\mathfrak{g}\to\mathfrak{u}^{+}$, the linear projection along $\mathfrak{l}+\mathfrak{u}^{-}$. Denote by $W$ the pre-image $\pi_{+}^{-1}(X)$ in $\mathcal{O}$, which consists of points $v\in\mathcal{O}$ such that $\lim_{\lambda\to 0}\phi_{\lambda}(v)\in X$.
We now show that $W$ is isomorphic to $T^{*}X$ following the argument of [Fu03, Lemma 3.7]. For any $x\in X$, let $T_{x}^{p}\coloneqq\{v\in T_{x}\mathcal{O}\mid(\phi_{\lambda})_{*}(v)=\lambda^{p}v\}$. In particular, $T_{x}^{0}=T_{x}X$. Take any $v_{1}\in T_{x}^{p}$ and $v_{2}\in T_{x}^{q}$. Since $\phi_{\lambda}^{*}\omega=\lambda\omega$, we have
$\lambda\omega(v_{1},v_{2})=\phi_{\lambda}^{*}\omega(v_{1},v_{2})=\omega((\phi_{\lambda})_{*}v_{1},(\phi_{\lambda})_{*}v_{2})=\lambda^{p+q}\omega(v_{1},v_{2}).$
Hence, if $p+q\neq 1$, then $\omega(v_{1},v_{2})=0$. This shows that $T_{x}^{0}\simeq T_{x}^{1}$ and $T_{x}^{p}=0$ for $p\neq 0,1$. Moreover, the symplectic form $\omega$ identifies $T_{x}^{1}$ with the dual of $T_{x}^{0}\cong T_{x}X$, hence we have $W\cong T^{*}X$.
Denote by $V$ the complement of $W$ in $\mathcal{O}$, which consists of points $v\in\mathcal{O}$ such that $\pi_{+}(v)\in\overline{X}\setminus X$. By [LSS88, Theorem 4.7] and [LSS88, Key Lemma, p. 41], we have
$\dim(\mathcal{O}\setminus W)=\dim(V)\leq\dim(\mathcal{O})-2,$
which implies that $k[T^{*}X]=k[\mathcal{O}]$, concluding the proof. ∎
Remark 4.10.
Pick $\mathfrak{g}$, $X$, and $\mathcal{O}$ in Table LABEL:table.Abelian. Let $U(\mathfrak{g})$ be the universal enveloping algebra of $\mathfrak{g}$, and let $\scr{D}(X)$ be the algebra of differential operators on $X$. By [LSS88, Theorem 5.2] and [LS89, Theorem, p. 2], there exists a completely prime, maximal ideal $J$ (depending on $\mathcal{O}$) of $U(\mathfrak{g})$ and an isomorphism $\psi\colon U(\mathfrak{g})/J\rightarrow\scr{D}(X)$. The Poincaré–Birkhoff–Witt filtration on $U(\mathfrak{g})$ gives a graded ring:
$R\coloneqq\operatorname{gr}U(\mathfrak{g})/\operatorname{gr}J={\sf S}^{\bullet}\mathfrak{g}/\operatorname{gr}J.$
Then the induced closed embedding $\operatorname{Spec}(R)\subset\mathfrak{g}^{*}$ is nothing but $\overline{\mathcal{O}}\subset\mathfrak{g}$ under the natural isomorphism $\mathfrak{g}\cong\mathfrak{g}^{*}$ induced by $\kappa$. On the other hand, the natural order filtration on $\scr{D}(X)$ induces a graded morphism $\operatorname{gr}\scr{D}(X)\rightarrow S(X)=k[T^{*}X]$. Then Theorem 4.9 implies that there exists a natural isomorphism $R\rightarrow S(X)$. This is a surprise to us because $\psi$ does not preserve the filtrations on each side (cf. [LS89, IV, § 1.9]).
Proof of Theorem 1.5.
Note that $\mathcal{O}_{1}=\mathcal{O}_{\min}$ and $X=\mathcal{O}_{\min}\cap\mathfrak{u}$, so the result follows from Theorem 4.9. ∎
Remark 4.11.
In Theorem 1.5, if $G/P$ is a projective space, i.e., $(\mathfrak{g},\mathfrak{p})=(\mathfrak{sl}(n+1),\mathfrak{p}_{1})$, the argument in Theorem 4.9 remains valid. However, as $\dim(\mathcal{O}\setminus W)=\dim(\mathcal{O})-1$ at the final step ([LSS88, Proposition 4.8]), we cannot conclude $k[T^{*}X]=k[\mathcal{O}]$.
Example 4.12.
Consider the hyperquadric cone $\hat{Q}\subset\mathbb{A}^{n+1}$ over a smooth projective hyperquadric in $\mathbb{P}^{n}$. Let $X=\hat{Q}\setminus\{0\}$; then $\overline{T^{*}X}$ is a symplectic variety of dimension $2n$, which is isomorphic to the minimal nilpotent orbit closure $\overline{\mathcal{O}_{\min}}$ in $\mathfrak{so}_{n+3}$. For instance, the basic affine space $\operatorname{SL}_{3}/U_{B}$ is isomorphic in codimension two to a hyperquadric cone in $\mathbb{A}^{6}$, hence $\overline{T^{*}\operatorname{SL}_{3}/U_{B}}$ is isomorphic to the closure of the minimal nilpotent orbit in $\mathfrak{so}_{8}$, a result proven in [Jia21, Theorem 1.3] via another method (cf. [Jia25]).
5. Parabolic basic affine spaces
Let $G$ be a connected reductive group, and let $P$ be a parabolic subgroup of $G$. Denote by $B\subset P$ the Borel subgroup of $G$ contained in $P$, and by $U_{P}$ the unipotent radical of $P$. Let $\mathfrak{u}_{P}^{\perp}\subset\mathfrak{g}^{*}$ be the annihilator of $\mathfrak{u}_{P}$. Let $X_{P}\coloneqq G/U_{P}$ be the parabolic basic affine space, which is quasi-affine.
The following result may be useful when considering the parabolic Ginzburg–Kazhdan conjecture for arbitrary $G$, and it shows that the problem can be further reduced to the case where $G$ is (almost) simple. Its proof is very similar to that of Corollary 1.3; thus, we omit it here for the interested reader.
Lemma 5.1.
Let $p\colon G_{1}\times\cdots\times G_{m}\times C\rightarrow G$ be an isogeny such that $G_{i}$ are semi-simple groups and $C$ is a torus. Let $P_{i}$ be the parabolic subgroup of $G_{i}$ such that $p(P_{1}\times\cdots\times P_{m}\times C)=P$.
  1. (5.1.1)
    The algebra $S(X_{P})$ is finitely generated if and only if $S(X_{P_{i}})$ is finitely generated for every $i$.
  2. (5.1.2)
    The affine closure $\overline{T^{*}X_{P}}$ is a $\mathbb{Q}$-factorial symplectic variety if and only if $\overline{T^{*}X_{P_{i}}}$ is symplectic for every $i$.
As an immediate consequence, one can derive from [GW25, Corollary 1.9] the following results:
Corollary 5.2.
Let $G$ be a reductive group and let $P$ be a parabolic subgroup of $G$. If the minimal connected normal algebraic subgroups of $[G,G]$ are of type A, then $S(X_{P})$ is finitely generated and $\overline{T^{*}X_{P}}$ is symplectic.
Proof.
Let $G_{i}$’s be the minimal connected algebraic subgroups of $[G,G]$, $1\leq i\leq r$. Then $G$ is an almost-direct product $G_{1}\times\cdots\times G_{r}\times C\rightarrow G$, where $C$ is a torus. Moreover, after passing to the universal covering of $G_{i}$, we may assume $G_{i}=\operatorname{SL}_{n_{i}}$ for some $n_{i}\geq 2$ by assumption. Then the statement follows from Lemma 5.1 and [GW25, Corollary 1.9]. ∎
From now on, after passing to an isogeny of simply connected type, we shall always assume that $G$ is connected and semi-simple. Note that $T^{*}X_{P}$ can be naturally identified with the geometric quotient
$(G\times\mathfrak{u}_{P}^{\perp})/U_{P}\coloneqq G\times_{U_{P}}\mathfrak{u}_{P}^{\perp},$
where $U_{P}$ acts on $G$ by right multiplication and on $\mathfrak{u}_{P}^{\perp}$ by the coadjoint action. The ring $S(X_{P})$ is isomorphic to the $U_{P}$-invariant subring $k[G\times\mathfrak{u}_{P}^{\perp}]^{U_{P}}$, and $\overline{T^{*}X_{P}}$ is isomorphic to the invariant-theoretic quotient $(G\times\mathfrak{u}_{P}^{\perp})/\!\!/U_{P}$. Let $M$ be the homogeneous bundle $G\times_{U_{B}}\mathfrak{u}_{P}^{\perp}$ over $G/U_{B}$. Then one obtains:
Lemma 5.3.
The ring $S(X_{P})$ is finitely generated (resp. has rational singularities) if and only if $k[M]$ is so.
Proof.
Let $P=U_{P}L_{P}$ be the Levi decomposition. Since $P$ is contained in the normalizer of $U_{P}$ in $G$, the natural $U_{P}$-action on $G\times\mathfrak{u}_{P}^{\perp}$ extends to a $P$-action, and thus the reductive group $L_{P}$ acts naturally on $G\times\mathfrak{u}_{P}^{\perp}$. Denote by $U_{B}$ the unipotent radical of $B$ and by $U_{L}$ a unipotent subgroup such that $U_{B}=U_{P}\rtimes U_{L}$. Then we have
$S(X_{P})^{U_{L}}\cong(k[G\times\mathfrak{u}_{P}^{\perp}]^{U_{P}})^{U_{L}}=k[G\times\mathfrak{u}_{P}^{\perp}]^{U_{B}},$
where the $U_{L}$-action on $S(X_{P})$ is induced by the natural $L_{P}$-action on $T^{*}X_{P}$. Notice that $(G\times\mathfrak{u}_{P}^{\perp})/\!\!/U_{B}$ is just the affine closure $\overline{M}$ of $M$. Now applying Theorem 2.10 to $L_{P}$ and $S(X_{P})$ yields the desired result. ∎
In the following, we aim to give an interpretation of the problem via birational geometry. Let $E=G\times_{B}\mathfrak{u}_{P}^{\perp}$.
Lemma 5.4.
There exists a graded isomorphism $\mathcal{R}(\mathbb{P}E^{*})\cong\mathcal{R}(\mathbb{P}M^{*})$.
Proof.
Let $\mathbb{T}$ be the torus $B/U_{B}$. Then the following morphism
$M=G\times_{U_{B}}\mathfrak{u}_{P}^{\perp}\longrightarrow G\times_{B}\mathfrak{u}_{P}^{\perp}=:E$
is exactly the geometric quotient of $M$ by the natural induced $\mathbb{T}$-action on $M$. Let $\mathbb{P}M^{*}$ and $\mathbb{P}E^{*}$ be the projectivizations of dual bundles. The induced morphism $q\colon\mathbb{P}M^{*}\rightarrow\mathbb{P}E^{*}$ is exactly the geometric quotient by the induced $\mathbb{T}$-action on $\mathbb{P}M^{*}$. The statement can then be derived from Theorem 2.3 immediately. ∎
Proof of Theorem 1.6.
By Lemma 5.1, we may assume that $G$ is simply connected. As $\mathbb{X}(U_{B})=1$, the divisor class group $\operatorname{Cl}(G/U_{B})$ is trivial by Lemma 4.2, and we have the following isomorphism
$\bigoplus_{p\geq 0}H^{0}(\mathbb{P}M^{*},\scr{O}_{\mathbb{P}M^{*}}(p))$ $\cong\mathcal{R}(\mathbb{P}M^{*})$ since $\text{since }\operatorname{Cl}(\mathbb{P}M^{*})\cong\mathbb{Z}\scr{O}_{\mathbb{P}M^{*}}(1)$
[…]
Now the claims follow from Theorem 2.7, Proposition 3.2, Lemma 5.3, and the following natural isomorphism
$k[M]\cong\bigoplus_{p\geq 0}H^{0}(G/U_{B},{\sf S}^{p}M^{*})\cong\bigoplus_{p\geq 0}H^{0}(\mathbb{P}M^{*},\scr{O}_{\mathbb{P}M^{*}}(p)).$
This finishes the proof. ∎
Let $F=G\times_{P}\mathfrak{u}_{P}^{\perp}$. Then we have a natural fibration $\mathbb{P}E^{*}\rightarrow\mathbb{P}F^{*}$. In particular, if $\mathbb{P}E^{*}$ is of Fano type, then $\mathbb{P}F^{*}$ is so by Theorem 2.7. As evidence for Conjecture 1.7, we confirm this corollary in the following.
Proposition 5.5.
The projectivization $\mathbb{P}F^{*}$ is weak Fano.
Proof.
Consider the following exact sequence of $P$-modules
$0\longrightarrow\mathfrak{u}_{P}\longrightarrow\mathfrak{g}\longrightarrow\mathfrak{g}/\mathfrak{u}_{P}\longrightarrow 0$
Note that $\mathfrak{g}/\mathfrak{u}_{P}\cong(\mathfrak{u}_{P}^{\perp})^{*}$ and $\mathfrak{u}_{P}\cong(\mathfrak{g}/\mathfrak{p})^{*}$, hence we have the following exact sequence of vector bundles over $G/P$:
$0\to T^{*}_{G/P}\to G/P\times\mathfrak{g}\to F^{*}\to 0$ (5.5.1)
Being a quotient of a trivial bundle, the vector bundle $F^{*}$ is globally generated. Moreover, we have $c_{1}(F^{*})=-c_{1}(T^{*}(G/P))=-K_{G/P}$.
Let $p\colon\mathbb{P}F^{*}\to G/P$ be the natural projection. Using the relative Euler sequence, we have $c_{1}(T_{p})=p^{*}c_{1}(F)+r\zeta$, where $r$ is the rank of $F$ and $\zeta=c_{1}(\mathcal{O}_{\mathbb{P}F^{*}}(1))$.
$-K_{\mathbb{P}F^{*}}=c_{1}(T_{p})-p^{*}K_{G/P}=p^{*}c_{1}(F)+r\zeta-p^{*}K_{G/P}=r\zeta.$
Hence $\mathbb{P}F^{*}$ is weak Fano if and only if $\zeta$ is nef and big. Moreover, as $F^{*}$ is globally generated, the divisor $\zeta$ is a nef divisor. Hence, to show $\zeta$ is big, we only need to compute its top self-intersection, or equivalently, the top Segre class of $F$.
By [EH16, Proposition 10.3] and the exact sequence (5.5.1), we have
$s(F)=\frac{1}{c(F)}=c(T_{G/P}),$
where $s(F)$ is the total Segre class of $F$ and $c(T_{G/P})$ is the total Chern class of $T_{G/P}$. It follows that $s_{\text{top}}(F)=c_{\text{top}}(G/P)$, which is just the topological Euler number of $G/P$, hence strictly positive. ∎
Example 5.6.
Let $G=\mathrm{SL}_{r+1}$ and let $P$ be the parabolic subgroup which is the stabilizer of an $r$-dimensional linear subspace in $k^{r+1}$. Then, according to [AT05, Example 3], we have:
$G/U_{P}\cong\{f\in\mathrm{Hom}(k^{r},k^{r+1})\mid f\text{ is injective}\}\subset\mathrm{Hom}(k^{r},k^{r+1}).$
Note that the complement has codimension two, hence $\overline{G/U_{P}}=\mathrm{Hom}(k^{r},k^{r+1})$. It follows that $T^{*}(G/U_{P})\subset T^{*}\mathrm{Hom}(k^{r},k^{r+1})$ has a complement of codimension two, from which we deduce that $\overline{T^{*}(G/U_{P})}=T^{*}\mathrm{Hom}(k^{r},k^{r+1})$.
Example 5.7.
Let $G=\mathrm{Sp}(4)$ and $P$ the stabilizer of a Lagrangian plane in $k^{4}$. By [AT05, Example 4], the affine closure of $G/U_{P}$ is isomorphic to a hyperquadric cone $\hat{Q}\subset\mathbb{A}^{8}$. As $G/U_{P}\subset\hat{Q}$ has a small boundary, we have $k[T^{*}(G/U_{P})]=k[T^{*}(\hat{Q}\setminus\{0\})]$, which implies that $\overline{T^{*}(G/U_{P})}$ is isomorphic to the minimal nilpotent orbit closure in $\mathfrak{so}_{10}$ by Example 4.12.